# Beyond Reference-Based Metrics: Analyzing Behaviors of Open LLMs on Data-to-Text Generation

### v1

@@ Appendix F Full Results
The Gigabyte GSmart Mika M3 is a smartphone that boasts a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels. It runs on Android 4.4.2 (KitKat) and is powered by a quad-core 1.3 GHz Cortex-A7 processor with a Mali-400MP2 GPU. The device comes with 8GB of internal storage and 1GB of RAM, which can be expanded via a dedicated microSDHC card slot. The primary camera is a 13-megapixel autofocus lens with an LED flash, while the secondary camera is an 8-megapixel lens. Connectivity options include Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The smartphone is equipped with a 1900mAh removable Li-Ion battery that provides up to 4 hours of talk time and up to 150 hours of standby time. The device is available in Navy Blue and supports dual SIM cards (Mini-SIM/ Micro-SIM).
The Gigabyte GSmart Mika M3 is a smartphone that boasts a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels. It runs on Android 4.4.2 (KitKat) and is powered by a quad-core 1.3 GHz Cortex-A7 processor with a Mali-400MP2 GPU. The device comes with 8GB of internal storage and 1GB of RAM, which can be expanded via a dedicated microSDHC card slot. The primary camera is a 13-megapixel autofocus lens with an LED flash, while the secondary camera is an 8-megapixel lens. Connectivity options include Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The smartphone is equipped with a 1900mAh removable Li-Ion battery that provides up to 4 hours of talk time and up to 150 hours of standby time. The device is available in Navy Blue and supports dual SIM cards (Mini-SIM/ Micro-SIM).
GPT-3.5
@@
The GSmart Mika M3, also known as the Gigabyte GSmart Mika M3, is a smartphone that was announced in October 2014 and released in November of the same year. It features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels and is powered by a Quad-core 1.3 GHz Cortex-A7 processor. The phone comes with 1GB of RAM and 8GB of internal storage, which can be expanded via a microSDHC card slot. The main camera is 13MP with LED flash, while the selfie camera is 8MP. The phone runs on Android 4.4.2 (KitKat) and is equipped with a 1900mAh removable battery. Other features include dual SIM capability, Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The GSmart Mika M3 is available in the color Navy Blue and has a sleek and compact design, measuring 144 x 70.5 x 8.3 mm and weighing 146g.
The GSmart Mika M3, also known as the Gigabyte GSmart Mika M3, is a smartphone that was announced in October 2014 and released in November of the same year. It features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels and is powered by a Quad-core 1.3 GHz Cortex-A7 processor. The phone comes with 1GB of RAM and 8GB of internal storage, which can be expanded via a microSDHC card slot. The main camera is 13MP with LED flash, while the selfie camera is 8MP. The phone runs on Android 4.4.2 (KitKat) and is equipped with a 1900mAh removable battery. Other features include dual SIM capability, Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The GSmart Mika M3 is available in the color Navy Blue and has a sleek and compact design, measuring 144 x 70.5 x 8.3 mm and weighing 146g.
Table 8: An example of the model outputs on gsmarena (see the input in Figure 8).
@@
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
@@
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk at their home arena. Kristall Saratov won the game with a final score of 1-5. Kristall Saratov’s home team scored one goal in the first period, while HC Norilsk scored two goals in the second and third periods. The game lasted for 60 minutes, with Kristall Saratov’s home team leading the game for the entire duration.
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk at their home arena. Kristall Saratov won the game with a final score of 1-5. Kristall Saratov’s home team scored one goal in the first period, while HC Norilsk scored two goals in the second and third periods. The game lasted for 60 minutes, with Kristall Saratov’s home team leading the game for the entire duration.
Mistral
@@
On November 28th, 2023, the VHL 23/24 season game between Kristall Saratov and HC Norilsk took place. The game was played at the home arena of Kristall Saratov, with the home team winning 1-0. The game was played in three periods, with the first period ending in a 0-0 tie, the second period ending in a 2-3 score in favor of HC Norilsk, and the third period ending in a 1-0 score in favor of Kristall Saratov. The game ended in normal time, with a total score of 3-1 in favor of HC Norilsk.
On November 28th, 2023, the VHL 23/24 season game between Kristall Saratov and HC Norilsk took place. The game was played at the home arena of Kristall Saratov, with the home team winning 1-0. The game was played in three periods, with the first period ending in a 0-0 tie, the second period ending in a 2-3 score in favor of HC Norilsk, and the third period ending in a 1-0 score in favor of Kristall Saratov. The game ended in normal time, with a total score of 3-1 in favor of HC Norilsk.
Zephyr
@@
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk. The first period saw two quick goals from Norilsk, with one in the second and third periods as well, resulting in a dominant 5-0 victory. Kristall Saratov managed to score a single goal in the third period, but it was not enough to overcome Norilsk’s lead. The final score was 1-5 in favor of HC Norilsk.
In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk. The first period saw two quick goals from Norilsk, with one in the second and third periods as well, resulting in a dominant 5-0 victory. Kristall Saratov managed to score a single goal in the third period, but it was not enough to overcome Norilsk’s lead. The final score was 1-5 in favor of HC Norilsk.
GPT-3.5
@@
In a match from the VHL 23/24 season, Kristall Saratov faced off against HC Norilsk. The game took place on November 28, 2023, at 15:30. Despite a goal from Kristall Saratov in the third period, HC Norilsk emerged victorious with a final score of 5-1. The away team dominated the game, scoring 2 goals in the first period and 3 goals in the second period. Kristall Saratov struggled to keep up, resulting in a clear win for HC Norilsk.
In a match from the VHL 23/24 season, Kristall Saratov faced off against HC Norilsk. The game took place on November 28, 2023, at 15:30. Despite a goal from Kristall Saratov in the third period, HC Norilsk emerged victorious with a final score of 5-1. The away team dominated the game, scoring 2 goals in the first period and 3 goals in the second period. Kristall Saratov struggled to keep up, resulting in a clear win for HC Norilsk.
Table 9: An example of the model outputs on ice_hockey (see the input in Figure 9).
@@
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
Life expectancy at birth in Djibouti has increased significantly over the past few decades, with a notable increase from 40.5665 years in 1950 to 62.3049 years in 2021.
Life expectancy at birth in Djibouti has increased significantly over the past few decades, with a notable increase from 40.5665 years in 1950 to 62.3049 years in 2021.
Mistral
This chart displays the life expectancy at birth in Djibouti from 1950 to 2021. The data shows a steady increase in life expectancy over the years, with a significant jump in the 1970s and 1980s. In 2021, the average life expectancy at birth in Djibouti was 62.3 years.
This chart displays the life expectancy at birth in Djibouti from 1950 to 2021. The data shows a steady increase in life expectancy over the years, with a significant jump in the 1970s and 1980s. In 2021, the average life expectancy at birth in Djibouti was 62.3 years.
Zephyr
This line graph illustrates the evolution of life expectancy at birth in Djibouti from 1950 to 2021. Initially, the average number of years that a newborn could expect to live was around 40.5 years in 1950, and it gradually increased over the following decades, reaching 62.3 years in 2021. The trend shows a steady improvement, with some fluctuations in the 1980s and 1990s. The data presented here is based on the sex- and age-specific death rates prevailing at the time of birth for a specific year in a given country, territory, or geographic area.
This line graph illustrates the evolution of life expectancy at birth in Djibouti from 1950 to 2021. Initially, the average number of years that a newborn could expect to live was around 40.5 years in 1950, and it gradually increased over the following decades, reaching 62.3 years in 2021. The trend shows a steady improvement, with some fluctuations in the 1980s and 1990s. The data presented here is based on the sex- and age-specific death rates prevailing at the time of birth for a specific year in a given country, territory, or geographic area.
GPT-3.5
The chart shows the life expectancy at birth in Djibouti from 1950 to 2021. The life expectancy has steadily increased over the years, from 40.57 years in 1950 to 62.30 years in 2021, with some fluctuations in between. The data reflects the average number of years a newborn could expect to live if exposed to the prevailing death rates at the time of birth.
The chart shows the life expectancy at birth in Djibouti from 1950 to 2021. The life expectancy has steadily increased over the years, from 40.57 years in 1950 to 62.30 years in 2021, with some fluctuations in between. The data reflects the average number of years a newborn could expect to live if exposed to the prevailing death rates at the time of birth.
Table 10: An example of the model outputs on owid (see the input in Figure 10).
@@
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
Driven Brands is an American franchise holding company that owns, operates, and franchises automotive services businesses in the United States. The company was founded in 1996 and is headquartered in Charlotte, North Carolina. It operates several brands, including Maaco, Meineke, Econo Lube, and CARSTAR.
Driven Brands is an American franchise holding company that owns, operates, and franchises automotive services businesses in the United States. The company was founded in 1996 and is headquartered in Charlotte, North Carolina. It operates several brands, including Maaco, Meineke, Econo Lube, and CARSTAR.
Mistral
Driven Brands is a business operating in the automotive services industry in the United States of America.
Driven Brands is a business operating in the automotive services industry in the United States of America.
Zephyr
Driven Brands is a business operating in the automotive services industry based in the United States of America. As an instance of the broader category of businesses, Driven Brands provides goods or services to customers in exchange for financial remuneration.
Driven Brands is a business operating in the automotive services industry based in the United States of America. As an instance of the broader category of businesses, Driven Brands provides goods or services to customers in exchange for financial remuneration.
GPT-3.5
Driven Brands is a business based in the United States of America, operating within the automotive services industry.
Driven Brands is a business based in the United States of America, operating within the automotive services industry.
Table 11: An example of the model outputs on wikidata (see the input in Figure 11).
Table 12: The average numbers of errors per output for each domain (lower is better). We also include the average number of tokens per output in the rightmost column. See Table 5 for aggregated results.
@@
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ Tok.
Llama2 openweather 5.77 4.02 1.94 2.06 0.34 0.52 0.43 0.16 8.48 6.76 166.7
gsmarena 1.88 0.65 0.37 0.89 0.33 0.19 0.11 0.09 2.69 1.82 90.3
ice_hockey 1.82 1.45 0.85 0.96 0.04 0.09 0.06 0.12 2.77 2.62 63.3
owid 2.18 1.25 0.22 0.58 0.22 0.28 0.01 0.07 2.63 2.18 60.3
wikidata 0.65 0.15 1.13 1.58 0.08 0.09 0.06 0.07 1.92 1.89 38.5
@@
gsmarena 1.86 0.51 0.80 1.10 0.54 0.29 0.05 0.12 3.25 2.02 146.3
ice_hockey 1.91 1.44 0.72 0.89 0.13 0.14 0.05 0.13 2.81 2.60 92.4
owid 2.67 2.39 0.18 0.40 0.41 0.29 0.04 0.07 3.30 3.15 91.1
wikidata 0.40 0.10 0.32 0.71 0.27 0.13 0.32 0.38 1.31 1.32 51.0
@@
gsmarena 1.72 0.34 0.93 1.04 0.85 0.23 0.03 0.10 3.53 1.71 142.8
ice_hockey 1.49 0.88 0.19 0.60 0.08 0.10 0.02 0.07 1.78 1.65 83.1
owid 2.51 1.63 0.20 0.52 0.43 0.22 0.01 0.10 3.15 2.47 85.2
wikidata 0.33 0.09 0.24 0.70 0.23 0.11 0.21 0.36 1.01 1.26 48.1
@@
gsmarena 1.30 0.20 1.13 0.80 0.63 0.21 0.03 0.14 3.09 1.35 129.5
ice_hockey 0.72 0.81 0.12 0.48 0.05 0.07 0.00 0.08 0.89 1.44 84.4
owid 1.81 0.63 0.03 0.25 0.35 0.17 0.00 0.01 2.19 1.06 62.2
wikidata 0.14 0.05 0.13 0.35 0.11 0.06 0.06 0.05 0.44 0.51 35.7
@@
$\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$
Llama2 openweather 0.96 0.74 0.70 0.70 0.26 0.29 0.25 0.12 1.00 0.88
gsmarena 0.83 0.38 0.28 0.48 0.21 0.18 0.06 0.05 0.90 0.74
ice_hockey 0.77 0.69 0.49 0.50 0.04 0.07 0.05 0.10 0.93 0.88
owid 0.92 0.68 0.18 0.37 0.19 0.24 0.01 0.06 0.93 0.87
wikidata 0.42 0.12 0.63 0.79 0.07 0.09 0.06 0.05 0.83 0.87
[…]
gsmarena 0.89 0.33 0.45 0.53 0.33 0.18 0.05 0.11 0.99 0.72
ice_hockey 0.83 0.73 0.52 0.60 0.13 0.12 0.04 0.09 0.96 0.87
owid 0.91 0.74 0.17 0.27 0.29 0.24 0.04 0.06 0.98 0.85
wikidata 0.29 0.08 0.25 0.43 0.19 0.10 0.20 0.24 0.70 0.66
[…]
gsmarena 0.87 0.23 0.53 0.47 0.45 0.19 0.02 0.08 0.99 0.60
ice_hockey 0.79 0.54 0.16 0.38 0.08 0.10 0.02 0.06 0.84 0.75
owid 0.93 0.68 0.14 0.31 0.30 0.20 0.01 0.07 0.96 0.83
wikidata 0.28 0.06 0.20 0.48 0.21 0.07 0.15 0.22 0.60 0.64
[…]
gsmarena 0.69 0.15 0.68 0.42 0.34 0.13 0.03 0.11 0.93 0.56
ice_hockey 0.52 0.64 0.09 0.30 0.05 0.07 0.00 0.07 0.56 0.76
owid 0.90 0.42 0.02 0.13 0.30 0.14 0.00 0.01 0.92 0.57
wikidata 0.12 0.04 0.11 0.26 0.10 0.06 0.06 0.05 0.33 0.37
Table 13: The ratio of outputs containing at least one error for each domain (lower is better). See Table 5 for aggregated results.
Table 12: The average numbers of errors per output for each domain (lower is better). We also include the average number of tokens per output in the rightmost column. See Table 5 for aggregated results.
Table 3: The average numbers of errors per output (lower is better) based on GPT-4 ($\mathcal{E}_{\text{gpt}}$) and human annotators ($\mathcal{E}_{\text{hum}}$). We also include the average number of tokens per output in the rightmost column. The results of the best open LLM are emphasized.
Figure 5: The instructions given to the human annotators.

### v3

@@ Appendix F Full Results
Zephyr The Gigabyte GSmart Mika M3 is a smartphone that boasts a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels. It runs on Android 4.4.2 (KitKat) and is powered by a quad-core 1.3 GHz Cortex-A7 processor with a Mali-400MP2 GPU. The device comes with 8GB of internal storage and 1GB of RAM, which can be expanded via a dedicated microSDHC card slot. The primary camera is a 13-megapixel autofocus lensNC with an LED flash, while the secondary camera is an 8-megapixel lens. Connectivity options include Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The smartphone is equipped with a 1900mAh removable Li-Ion battery that provides up to 4 hoursI of talk time and up to 150 hours of standby time. The device is available in Navy Blue and supports dual SIM cards (Mini-SIM/ Micro-SIM). The Gigabyte GSmart Mika M3 is a smartphone that boasts a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels. It runs on Android 4.4.2 (KitKat) and is powered by a quad-core 1.3 GHz Cortex-A7 processor with a Mali-400MP2 GPU. The device comes with 8GB of internal storage and 1GB of RAM, which can be expanded via a dedicated microSDHC card slot. The primary camera is a 13-megapixel autofocus lens with an LED flash, while the secondary camera is an 8-megapixel lens. Connectivity options include Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The smartphone is equipped with a 1900mAh removable Li-Ion battery that provides up to 4 hours of talk timeI and up to 150 hours of standby time. The device is available in Navy Blue and supports dual SIM cards (Mini-SIM/ Micro-SIM).
@@
GPT-3.5 The GSmart Mika M3, also known as the Gigabyte GSmart Mika M3, is a smartphone that was announced in October 2014 and released in November of the same year. It features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels and is powered by a Quad-core 1.3 GHz Cortex-A7 processor. The phone comes with 1GB of RAM and 8GB of internal storage, which can be expanded via a microSDHC card slot. The main camera is 13MP with LED flash, while the selfie camera is 8MP. The phone runs on Android 4.4.2 (KitKat) and is equipped with a 1900mAh removable battery. Other features include dual SIM capability, Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The GSmart Mika M3 is available in the color Navy Blue and has a sleek and compact design, measuring 144 x 70.5 x 8.3 mm and weighing 146g. The GSmart Mika M3, also known as the Gigabyte GSmart Mika M3I, is a smartphone that was announced in October 2014 and released in November of the same year. It features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels and is powered by a Quad-core 1.3 GHz Cortex-A7 processor. The phone comes with 1GB of RAM and 8GB of internal storage, which can be expanded via a microSDHC card slot. The main camera is 13MP with LED flashNC, while the selfie camera is 8MPNC. The phone runs on Android 4.4.2 (KitKat) and is equipped with a 1900mAh removable batteryNC. Other features include dual SIM capability, Wi-Fi, Bluetooth, GPS, FM radio, and microUSB 2.0. The GSmart Mika M3 is available in the color Navy Blue and has a sleek and compact designNC, measuring 144 x 70.5 x 8.3 mm and weighing 146g.
Table 9: An example of the model outputs on gsmarena (see the input in Table 9).
@@
Model Human annotations ($\mathcal{E}_{\text{hum}}$) GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
@@
Llama2 In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk at their home arena. Kristall Saratov won the gameI with a final score of 1-5. Kristall Saratov’s home team scored one goal in the first periodI, while HC Norilsk scored two goals in the second and third periodsI. The game lasted for 60 minutesNC, with Kristall Saratov’s home team leading the game for the entire durationI. In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk at their home arena. Kristall Saratov won the game with a final score of 1-5.I Kristall Saratov’s home team scored one goal in the first periodI, while HC Norilsk scored two goals in the second and third periods.I The game lasted for 60 minutesNC, with Kristall Saratov’s home team leading the game for the entire duration.I
@@
Mistral On November 28th, 2023, the VHL 23/24 season game between Kristall Saratov and HC Norilsk took place. The game was played at the home arena of Kristall Saratov, with the home team winning 1-0I. The game was played in three periods, with the first period ending in a 0-0 tieI, the second period ending in a 2-3 score in favor of HC NorilskI, and the third period ending in a 1-0 score in favor of Kristall Saratov. The game ended in normal time, with a total score of 3-1 in favor of HC NorilskI. On November 28th, 2023, the VHL 23/24 season game between Kristall Saratov and HC Norilsk took place. The game was played at the home arena of Kristall Saratov, with the home team winning 1-0I. The game was played in three periods, with the first period ending in a 0-0 tie, the second period ending in a 2-3 score in favor of HC NorilskI, and the third period ending in a 1-0 score in favor of Kristall Saratov. The game ended in normal time, with a total score of 3-1 in favor of HC NorilskI.
@@
Zephyr In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk. The first period saw two quickNC goals from Norilsk, with one in the second and third periods as wellI, resulting in a dominant 5-0I victory. Kristall Saratov managed to score a single goal in the third period, but it was not enough to overcome Norilsk’s lead. The final score was 1-5 in favor of HC Norilsk. In the VHL 23/24 season, Kristall Saratov hosted HC Norilsk. The first period saw two quick goals from Norilsk, with one in the secondI and third periods as well, resulting in a dominant 5-0 victoryI. Kristall Saratov managed to score a single goal in the third period, but it was not enough to overcome Norilsk’s lead. The final score was 1-5 in favor of HC Norilsk.
@@
GPT-3.5 In a match from the VHL 23/24 season, Kristall Saratov faced off against HC Norilsk. The game took place on November 28, 2023, at 15:30I. Despite a goal from Kristall Saratov in the third period, HC Norilsk emerged victorious with a final score of 5-1. The away team dominated the game, scoring 2 goals in the first period and 3 goals in the second period. Kristall Saratov struggled to keep up, resulting in a clear win for HC Norilsk. In a match from the VHL 23/24 season, Kristall Saratov faced off against HC Norilsk. The game took place on November 28, 2023, at 15:30. Despite a goal from Kristall Saratov in the third period, HC Norilsk emerged victorious with a final score of 5-1. The away team dominated the game, scoring 2 goals in the first period and 3 goals in the second period. Kristall Saratov struggled to keep up, resulting in a clear win for HC Norilsk.
Table 10: An example of the model outputs on ice_hockey (see the input in Table 10).
@@
Model Human annotations ($\mathcal{E}_{\text{hum}}$) GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2 Life expectancy at birth in Djibouti has increased significantly over the past few decades, with a notable increase from 40.5665 years in 1950 to 62.3049 years in 2021. Life expectancy at birth in Djibouti has increased significantly over the past few decades, with a notable increase from 40.5665 years in 1950 to 62.3049 years in 2021.
Mistral This chart displays the life expectancy at birth in Djibouti from 1950 to 2021. The data shows a steady increase in life expectancy over the years, with a significant jump in the 1970s and 1980sI. In 2021, the average life expectancy at birth in Djibouti was 62.3 years. This chart displays the life expectancy at birth in Djibouti from 1950 to 2021. The data shows a steady increaseI in life expectancy over the years, with a significant jump in the 1970s and 1980sI. In 2021, the average life expectancy at birth in Djibouti was 62.3 yearsI.
Zephyr This line graph illustrates the evolution of life expectancy at birth in Djibouti from 1950 to 2021. Initially, the average number of years that a newborn could expect to live was around 40.5 years in 1950, and it gradually increased over the following decades, reaching 62.3 years in 2021. The trend shows a steady improvement, with some fluctuations in the 1980sI and 1990s. The data presented here is based on the sex- and age-specific death rates prevailing at the time of birth for a specific year in a given country, territory, or geographic area. This line graph illustrates the evolution of life expectancy at birth in Djibouti from 1950 to 2021. Initially, the average number of years that a newborn could expect to live was around 40.5 years in 1950I, and it gradually increased over the following decades, reaching 62.3 years in 2021I. The trend shows a steady improvementM, with some fluctuations in the 1980s and 1990s. The data presented here is based on the sex- and age-specific death rates prevailing at the time of birth for a specific year in a given country, territory, or geographic area.
GPT-3.5 The chart shows the life expectancy at birth in Djibouti from 1950 to 2021. The life expectancy has steadily increased over the years, from 40.57 years in 1950 to 62.30 years in 2021, with some fluctuations in between. The data reflects the average number of years a newborn could expect to live if exposed to the prevailing death rates at the time of birth. The chart shows the life expectancy at birth in Djibouti from 1950 to 2021. The life expectancy has steadily increased over the yearsI, from 40.57 years in 1950I to 62.30 years in 2021I, with some fluctuations in between. The data reflects the average number of years a newborn could expect to live if exposed to the prevailing death rates at the time of birth.
Table 11: An example of the model outputs on owid (see the input in Table 11).
@@
Model Human annotations ($\mathcal{E}_{\text{hum}}$) GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2 Driven Brands is an American franchise holdingNC company that owns, operates, and franchisesNC automotive services businesses in the United States. The company was founded in 1996 and is headquartered in Charlotte, North Carolina. It operates several brands, including Maaco, Meineke, Econo Lube, and CARSTARNC. Driven Brands is an American franchise holding company that owns, operates, and franchises automotive services businesses in the United States. The company was founded in 1996NC and is headquartered in Charlotte, North CarolinaNC. It operates several brands, including Maaco, Meineke, Econo Lube, and CARSTAR.
Mistral Driven Brands is a business operating in the automotive services industry in the United States of America. Driven Brands is a business operating in the automotive services industry in the United States of America.
Zephyr Driven Brands is a business operating in the automotive services industry based in the United States of America. As an instance of the broader category of businesses, Driven Brands provides goods or services to customers in exchange for financial remunerationNC. Driven Brands is a business operating in the automotive services industry based in the United States of America. As an instance of the broader category of businesses,O Driven Brands provides goods or servicesI to customers in exchange for financial remuneration.
GPT-3.5 Driven Brands is a business based in the United States of America, operating within the automotive services industry. Driven Brands is a business based in the United States of America, operating within the automotive services industry.
Table 12: An example of the model outputs on wikidata (see the input in Table 12).
@@
$\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ Tok.
Llama 2 openweather 4.27 6.99 2.23 1.93 0.58 0.20 0.16 0.26 7.24 9.38 166.7
gsmarena 0.67 1.96 0.87 0.31 0.20 0.24 0.09 0.10 1.83 2.61 90.3
ice_hockey 1.47 1.98 0.99 0.90 0.09 0.00 0.12 0.05 2.67 2.93 63.3
owid 1.30 2.26 0.58 0.15 0.32 0.13 0.07 0.00 2.27 2.54 60.3
wikidata 0.15 0.74 1.59 1.24 0.08 0.03 0.07 0.04 1.89 2.05 38.5
@@
gsmarena 0.51 2.02 1.10 0.94 0.30 0.25 0.14 0.02 2.05 3.23 146.3
ice_hockey 1.47 1.79 0.92 0.76 0.14 0.08 0.13 0.10 2.66 2.73 92.4
owid 2.42 3.11 0.40 0.18 0.29 0.27 0.07 0.02 3.18 3.58 91.1
wikidata 0.10 0.45 0.72 0.40 0.13 0.19 0.38 0.31 1.33 1.35 51.0
@@
gsmarena 0.33 2.24 1.04 1.00 0.23 0.48 0.11 0.01 1.71 3.73 142.8
ice_hockey 0.89 1.57 0.61 0.20 0.11 0.04 0.07 0.01 1.68 1.82 83.1
owid 1.68 2.64 0.49 0.16 0.22 0.30 0.10 0.01 2.49 3.11 85.2
wikidata 0.09 0.36 0.72 0.29 0.12 0.21 0.36 0.21 1.29 1.07 48.1
@@
gsmarena 0.20 1.64 0.80 1.42 0.21 0.34 0.17 0.01 1.38 3.41 129.5
ice_hockey 0.81 0.76 0.46 0.17 0.07 0.04 0.09 0.01 1.43 0.98 84.4
owid 0.64 1.87 0.25 0.02 0.17 0.27 0.01 0.00 1.07 2.16 62.2
wikidata 0.05 0.18 0.36 0.16 0.06 0.10 0.05 0.10 0.52 0.54 35.7
Table 13: The average numbers of errors per output for each domain (lower is better). We also include the average number of tokens per output in the rightmost column. See section 5 for aggregated results.
@@
$\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$ $\mathcal{E}_{\text{hum}}$ $\mathcal{E}_{\text{gpt}}$
Llama 2 openweather 76% 98% 71% 68% 30% 14% 12% 22% 90% 100%
gsmarena 38% 86% 49% 24% 18% 16% 5% 7% 74% 94%
ice_hockey 70% 77% 51% 55% 7% 0% 10% 5% 89% 94%
owid 70% 92% 37% 14% 24% 12% 6% 0% 88% 93%
wikidata 12% 47% 79% 63% 8% 2% 5% 4% 87% 89%
Mistral openweather 78% 99% 61% 16% 38% 18% 17% 4% 92% 100%
gsmarena 33% 89% 54% 54% 19% 17% 12% 2% 73% 99%
ice_hockey 74% 83% 62% 52% 12% 8% 9% 7% 88% 97%
owid 75% 96% 27% 8% 24% 25% 6% 2% 86% 97%
wikidata 8% 34% 44% 29% 10% 17% 24% 27% 67% 72%
Zephyr openweather 82% 99% 45% 25% 23% 25% 14% 0% 90% 100%
gsmarena 22% 91% 47% 52% 19% 30% 9% 1% 61% 99%
ice_hockey 54% 78% 39% 17% 11% 4% 6% 1% 76% 87%
owid 70% 93% 31% 9% 20% 26% 7% 1% 85% 95%
wikidata 6% 29% 49% 22% 8% 18% 22% 18% 66% 66%
GPT-3.5 openweather 64% 98% 32% 12% 23% 25% 5% 0% 75% 99%
gsmarena 15% 74% 42% 58% 13% 19% 12% 1% 57% 97%
ice_hockey 64% 50% 30% 12% 7% 4% 8% 1% 76% 57%
owid 43% 89% 13% 2% 14% 24% 1% 0% 57% 90%
wikidata 4% 14% 27% 14% 6% 9% 5% 9% 38% 36%
Table 14: The percentage of outputs containing at least one error for each domain (lower is better). See section 5 for aggregated results.